OpenStudy (anonymous):

Einstein's formula for relativity is E = mc2 where E = energy, m = mass, and c = the speed of light. Solve E = mc2 for c a) c= square root of quantity E over m b)c = E over m c) c = square root of quantity m over E d) c = m over E

OpenStudy (unklerhaukus):

E = mc^2 To solve the equation for c, • divide both sides by m, then • take the square root of both sides

OpenStudy (anonymous):

can you show me how to do it.. like a visual example?

OpenStudy (unklerhaukus):

if you instead had F = kx^2 and were solving for x F/k = kx^2/k (dividing both sides by k) F/k = x^2 (simplifying the right hand side) √(F/k) = √x^2 (taking the square root ) √(F/k) = x (simplifying the right hand side) x= √(F/k) (answer)

OpenStudy (unklerhaukus):

Your first step is to divide both sides by m E = mc^2 . . .

OpenStudy (anonymous):

so A is the answer?

OpenStudy (anonymous):